# Diffusion Models for Reinforcement Learning: A Survey

### 11 Dec 2023 v2

@@ Abstract
Diffusion models have emerged as a prominent class of generative models, surpassing previous methods regarding sample quality and training stability. Recent works have shown the advantages of diffusion models in improving reinforcement learning (RL) solutions, including as trajectory planners, expressive policy classes, data synthesizers, etc. This survey aims to provide an overview of the advancements in this emerging field and hopes to inspire new avenues of research. First, we examine several challenges encountered by current RL algorithms. Then, we present a taxonomy of existing methods based on the roles played by diffusion models in RL and explore how the existing challenges are addressed. We further outline successful applications of diffusion models in various RL-related tasks while discussing the limitations of current approaches. Finally, we conclude the survey and offer insights into future research directions, focusing on enhancing model performance and applying diffusion models to broader tasks. We are actively maintaining a GitHub repository for papers and other related resources in applying diffusion models in RL  [^1] https://github.com/apexrl/Diff4RLSurvey.
@@ 1 Introduction
This survey centers its attention on the utilization of diffusion models in RL, with additional consideration given to methods incorporating diffusion models in the contexts of trajectory generation and imitation learning, primarily due to the evident interrelations between these fields. Section 2 elaborates on the aforementioned RL challenges, and discusses how diffusion models can help solve each challenge. Section 3 provides a background on the foundations of diffusion models and also covers two class of methods that are particularly important in RL-related applications: guided sampling and fast sampling. Section 4 illustrates what roles diffusion models play in RL among existing works. Section 5 discusses the contribution of diffusion models on different RL-related applications. In Section 6, we point out the limitations when applying diffusion models and compare them with the transformer-based methods. Section 7 summarizes the survey with a discussion on emerging new topics.
@@ 2.1 Restricted Expressiveness in Offline Learning
Online RL Sutton and Barto (2018); Arulkumaran et al. (2017) has been criticized for low sample efficiency, which makes it difficult to be applied in real-world scenarios. Offline RL Fujimoto et al. (2019); Kumar et al. (2020); Fujimoto and Gu (2021), which learns optimal policies purely from pre-collected datasets, obviates the need to interact with the environment during training, and can significantly improve sample efficiency. Directly applying off-policy RL methods Lillicrap et al. (2015); Haarnoja et al. (2018) to offline learning suffers from the extrapolation error problem Fujimoto et al. (2019). Existing works either penalize the value predictions on out-of-distribution samples Kumar et al. (2020); Nachum et al. (2019), or constrain the learning policy to be close to the data collection policy Wu et al. (2019); Kostrikov et al. (2021). However, penalties on the value function can make the learned policy over-conservative Lyu et al. (2022); for algorithms using policy constraints, since the policy is usually parameterized as a unimodal Gaussian, the restricted expressiveness makes it hard to fit the possibly diversified dataset. The reinforcement learning via supervised learning framework (RvS) Schmidhuber (2019); Srivastava et al. (2019) is now becoming another essential paradigm in offline RL, which bypasses the need for Q learning and is thus free of extrapolation errors. RvS learns a policy conditioned on the observed returns via supervised learning and then conditions the learned policy on a high enough return during online evaluations to generate desired behaviors Chen et al. (2021); Lee et al. (2022). Similar to policy constraining methods, RvS requires fitting the entire offline dataset. Therefore, the expressiveness of parametrized policies also matters in RvS. Since diffusion models can represent arbitrary normalizable distributions Neal and others (2011), they hold potential to effectively improve the performance of policy constraining and RvS algorithms on complex datasets.
@@ 2.2 Data Scarcity in Experience Replay
Off-policy and offline RL methods use different levels of experience replay to improve sample efficiency. Note that experience replay in some cases only refers to data reuse in off-policy RL. Here, we use the term to broadly refer to updating the current model with rollout data from other policies. In off-policy RL, although all previously collected experiences can be used for policy learning, the available data during training may still be inadequate for effective policy optimization due to the speed limit of simulations and the potentially large state and action spaces. In offline RL, as no further interactions are allowed, policy learning is more limited by the quality and coverage of the dataset. Inspired by the success of data augmentation in computer vision, some works implement similar augmentation techniques in RL to mitigate the data scarcity problem. RAD Laskin et al. (2020) uses typical image augmentation techniques such as random cropping or rotation to boost the learning efficiency in vision-based RL. Imre (2021) and Cho et al. (2022) use generative models, VAE Kingma and Welling (2013) and GAN Goodfellow et al. (2014), to augment the real dataset with synthetic data sampled from the learned data distribution. However, existing works either suffer from a lack of fidelity when using random augmentation or are limited to simple environments due to insufficient modeling ability of particular generative models, making it difficult for them to extend to more complex tasks. Diffusion models have already demonstrated notable performances surpassing previous generative models in domains including image and video synthesis Ho et al. (2020, 2022). When applied to RL data, diffusion models are more suitable for augmenting high-dimensional datasets with intricate interactions.
@@ 2.3 Compounding Error in Model-based Planning
Model-based RL (MBRL) Luo et al. (2022); Moerland et al. (2023) fits a model of dynamic transitions either from data obtained from online rollout or an offline dataset, and expects the model to facilitate decision-making. Common dynamic models mimic single-step state transitions and rewards in the dataset. Due to the limited data support and the possible stochasticity of the ground-truth transition function, there could be single-step errors when predicting with a neural network model. As a result, the compounding error problem arises when using the model for multistep planning Xiao et al. (2019), i.e., cumulative single-step errors can cause the planned states to deviate from the dataset distribution, further increasing the error in the subsequent model predictions. In contrast, diffusion models with powerful modeling ability of joint distributions can operate on the trajectory level and plan for all time steps simultaneously, offering better temporal consistency and less compounding errors.
@@ 2.4 Generalization in Multitask Learning
Normal RL algorithms lack generalization abilities at the task level Vithayathil Varghese and Mahmoud (2020). Even in the same environment, changing the reward function requires retraining a policy from scratch. Existing works studying online multitask RL Yu et al. (2020); Liu et al. (2021) attempt to learn the same policy on different task environments, suffering from the problem of conflicting gradients across multiple tasks, as well as low sample efficiency due to pure online learning. Recently, it has become a trending research direction to utilize a high-capacity model trained on multitask offline datasets and then deployed on new tasks with or without online fine-tuning Taiga et al. (2022); Oh et al. (2017). Transformer-based pre-training decision models like Gato Reed et al. (2022) achieve notable successes in multitask policy learning. However, they typically require high-quality datasets and come with large parameter sizes as well as corresponding high training and inference costs. How to design an algorithm that can efficiently fit multitask datasets of mixed quality and generalize to new tasks emerges as a vital issue in multitask RL. As a powerful class of generative models, diffusion models can handle multimodal distributions in multitask datasets, and adapt to new tasks by estimating the task distribution.
@@ 3 Foundations of Diffusion Models
This section provides the foundations of diffusion models. Two prominent formulations are presented: Denoising Diffusion Probabilistic Model (DDPM) Ho et al. (2020) and Score-based Generative Models Song et al. (2021). DDPM is widely used due to its simplicity, whereas the score-based formulation extends it to encompass continuous-time diffusion processes. Furthermore, guided sampling methods play a critical role in integrating diffusion models into RL frameworks. These methods can be divided into two main categories based on their approach to guiding the sampling process: classifier-guidance Dhariwal and Nichol (2021), which requires an extra classifier, and classifier-free guidance Ho and Salimans (2022), which makes guiding conditions as part of the model input. Additionally, in order to boost sampling speed especially during online interactions, fast sampling techniques have been adopted when using diffusion models in RL-related tasks Kang et al. (2023); Chi et al. (2023). Here we briefly cover some representative works in studying fasting sampling of diffusion models, including learning-based and learning-free methods.
@@ 3.1 Denoising Diffusion Probabilistic Model
Assuming that the real data, denoted as $x^{0}$, is sampled from an underlying distribution $q(x^{0})$, the Denoising Diffusion Probabilistic Model (DDPM) utilizes a parameterized diffusion process, represented as $p_{\theta}(x^{0})=\int p(x^{T})\prod_{t=1}^{T}p_{\theta}(x^{t-1}|x^{t})\mathop{}\!\mathrm{d}x^{1:T}$, to model how the pure noise, denoted as $x^{T}=\mathcal{N}(0,I)$, is denoised into real data $x^{0}$. In this formulation, each step of the diffusion process is represented by $x^{t}$, with $T$ indicating the total number of steps. It is important to note that both diffusion models and RL use time step notations; thus, following common practice in RL, we denote diffusion steps with superscripts and RL time steps with subscripts. The sequence $x^{T:0}$ is defined as a Markov chain with learned Gaussian transitions characterized by
$p_{\theta}(x^{t-1}|x^{t})=\mathcal{N}(\mu_{\theta}(x^{t},t),\Sigma(x^{t},t))~{}.$
If the process is reversed to $x^{0:T}$, each step is the forward transition $q(x^{t}|x^{t-1})$. This transition can be interpreted as adding Gaussian noise to the data according to a variance schedule $\beta^{1:T}$:
$x^{t}=\sqrt{\alpha^{t}}x^{t-1}+\sqrt{1-\alpha^{t}}\epsilon_{t}~{},$ (1)
where $\alpha^{t}=1-\beta^{t}$, $\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. $q$ also holds the Markov property as $q(x^{t}|x^{t-1})$ only depends on $x^{t-1}$. In general, $\alpha^{T}$ should satisfy $\lim_{T\rightarrow+\infty}\alpha^{T}\rightarrow 0$ to ensure that $x^{T}$ would be a pure standard Gaussian noise when $T$ is sufficiently large. And from Eq. (1), we can infer that
$x^{t}=\sqrt{\bar{\alpha}^{t}}x^{0}+\sqrt{1-\bar{\alpha}^{t}}{\epsilon(x^{t},t)}~{},$ (2)
where $\bar{\alpha}^{t}=\prod_{1}^{t}\alpha^{i}$ and ${\epsilon(x^{t},t)}\sim\mathcal{N}(0,\mathbf{I})$ is the integrated noise from step 0 to $t$. $\epsilon(x,t)$ is unknown and to be learned by a network. So far, we can calculate the distribution from step 0 to step $t$ by simply one calculation. As for the denoising process $q(x^{t-1}|x^{t})$, it is complex while $q(x^{t-1}|x^{t},x^{0})$ is relatively simpler. Based on Bayes Theorem and the Markov property of $q$, we have
$q(x^{t-1}|x^{0},x^{t})=\frac{q(x^{t-1}|x^{0})}{q(x^{t}|x^{0})}q(x^{t}|x^{t-1})=\mathcal{N}(\mu,\sigma^{2})~{},$
where
$\sigma^{2}(x^{t},t)$ $=\beta^{t}\frac{1-\bar{\alpha}^{t-1}}{1-\bar{\alpha}^{t}},$ (3)
$\mu(x^{t},t)$ $=\frac{1-\bar{\alpha}^{t-1}}{1-\bar{\alpha}^{t}}\sqrt{\alpha^{t}}x^{t}+\frac{\sqrt{\bar{\alpha}^{t-1}}\beta^{t}}{1-\bar{\alpha}^{t}}x^{0}~{}.$
From Eq. (2), we can eliminate $x^{0}$, yielding:
$\mu(x^{t},t)=\frac{1}{\sqrt{{\alpha}^{t}}}(x^{t}-\frac{\beta^{t}}{\sqrt{1-\bar{\alpha}^{t}}}\epsilon(x^{t},t))~{}.$ (4)
This allows us to sample $x^{T}$ from standard Gaussian noise and progressively denoise it step by step until we obtain $x^{0}$. However, the noise variable $\epsilon$ is still unknown. To address this, a network $\epsilon_{\theta}$, parameterized by $\theta$, is employed to learn the noise generation process. In general, the loss function is the distance between the generated $\mu(x^{t},t)$ using random variables $\epsilon$ and the network output $\epsilon_{\theta}$. Specially, $\mu(x^{t},t)$ is defined as in Eq. (4), and $\mu_{\theta}(x^{t},t)=\frac{1}{\sqrt{{\alpha}^{t}}}(x^{t}-\frac{\beta^{t}}{\sqrt{1-\bar{\alpha}^{t}}}\epsilon_{\theta}(x^{t},t))$.
Ho et al. (2020) discover that the following version of the loss function ignoring the weights has a better performance in experiments:
$\mathcal{L}_{t-1}\coloneqq\mathbb{E}_{x^{0},\epsilon}[\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}^{t}}x^{0}+(\sqrt{1-\bar{\alpha}^{t}})\epsilon,t)\|^{2}]~{}.$ (5)
Here, $\epsilon$ is sampled from the standard Gaussian distribution $\mathcal{N}(0,\mathbf{I})$ and $\mathcal{L}_{t-1}$ represents the loss function at diffusion step $t-1$, when $t>1$.
@@ 3.2 Score-based Generative Models
DDPM illustrates its iterative sampling process in a step-by-step manner, with each step taking place at a discrete time interval. Song et al. (2021) extend DDPM to continuous-time cases, where the sequence $x^{0},x^{1},\ldots,x^{T}$ is replaced with a continual function $x^{t},t\in[0,T]$. Now, the forwarding process can be described as a Stochastic Differential Equation (SDE):
@@
where $f(x,t)$ and $g(t)$ are pre-defined functions, and $\mathop{}\!\mathrm{d}w$ is the Brownian Motion. The example DDPM process in Eq. (1) is represented as $f(x,t)=-\frac{1}{2}\beta(t)x$, $g(t)=\sqrt{\beta(t)}~{}.$ According to Langevin Dynamics, the reverse of the forwarding process, as the sampling process, is described by a reverse-time SDE:
$\mathop{}\!\mathrm{d}x=[f(x,t)-g^{2}(t)\nabla_{x}\log q_{t}(x)]\mathop{}\!\mathrm{d}t+g(t)\mathop{}\!\mathrm{d}\bar{w}~{},$ (6)
where $\bar{w}$ is the reverse Brown Motion. The gradient term $\nabla_{x}\log q_{t}(x)$ is called the score function of the distribution $q_{t}(x)$. As the score function remains unknown, score matching Song et al. (2019) proposes that a parameterized score-based model $\epsilon_{\theta}(x,t)$ can be trained to approximate it. In practice, diffusion models usually match the scaled score function $-\sigma_{t}\nabla_{x}\log q_{t}(x)$. Thus, a time-dependent score model can be trained by minimizing the Fisher Divergence:
$\min_{\theta}\mathbb{E}_{x^{0}\sim q_{0},x^{t}\sim q_{0t}}\left[\|\epsilon_{\theta}(x^{t},t)+\sigma_{t}\nabla_{x^{t}}\log q_{0t}(x^{t}|x^{0})\|_{2}^{2}\right]~{}.$
Solving the SDE in Eq. (6) is difficult even with a trained score model. Alternatively, we can solve the corresponding Probability Flow ODE (PF-ODE), whose marginal distribution at any time is equal to that given by the original SDE in Eq. (6):
$\mathop{}\!\mathrm{d}x=\left[f(x,t)-\frac{1}{2}g^{2}(t)\nabla_{x}\log q_{t}(x)\right]\mathop{}\!\mathrm{d}t~{}.$ (7)
Given a trained score model, all terms in Eq. (7) are known thus it can be efficiently solved by various ODE solvers such as VODE Brown et al. (1989). In this case, a black-box ODE solver Dormand and Prince (1980) is highly recommended because it not only produces high-quality samples but also allows us to explicitly trade-off accuracy for efficiency.

### 23 Feb 2024 v4

@@ Abstract
Diffusion models surpass previous generative models in sample quality and training stability. Recent works have shown the advantages of diffusion models in improving reinforcement learning (RL) solutions. This survey aims to provide an overview of this emerging field and hopes to inspire new avenues of research. First, we examine several challenges encountered by RL algorithms. Then, we present a taxonomy of existing methods based on the roles of diffusion models in RL and explore how the preceding challenges are addressed. We further outline successful applications of diffusion models in various RL-related tasks. Finally, we conclude the survey and offer insights into future research directions. We are actively maintaining a GitHub repository for papers and other related resources in utilizing diffusion models in RL  [^2] https://github.com/apexrl/Diff4RLSurvey.
@@ 1 Introduction
This survey centers its attention on the utilization of diffusion models in RL, with additional consideration given to methods incorporating diffusion models in the contexts of trajectory generation and imitation learning, primarily due to the evident interrelations between these fields. Section 2 elaborates on the aforementioned RL challenges, and discusses how diffusion models can help solve each challenge. Section 3 provides a background on the foundations of diffusion models and also covers two class of methods that are particularly important in RL-related applications: guided sampling and fast sampling. Section 4 illustrates what roles diffusion models play in RL among existing works. Section 5 discusses the contribution of diffusion models on different RL-related applications. Section 6 summarizes the survey with a discussion on emerging new topics.
@@ 2.1 Restricted Expressiveness in Offline Learning
Due to the low-sample efficiency, online RL Sutton and Barto (2018) is challenging to be applied in real-world scenarios. Offline RL Fujimoto et al. (2019); Kumar et al. (2020) learns optimal policies from pre-collected datasets without environmental interaction, which can significantly improve sample efficiency. Directly applying off-policy RL to offline learning causes severe extrapolation errors Fujimoto et al. (2019). Existing works penalize value predictions on out-of-distribution samples Kumar et al. (2020) or limit the learning policy to be close to the data collecting policy Kostrikov et al. (2021). However, penalties imposed on the value function may result in an overly conservative policy Lyu et al. (2022); when imposing policy constraints on commonly used unimodal Gaussian parameterization, the restricted expressiveness makes it difficult to fit the possibly diverse dataset. Reinforcement learning via supervised learning framework (RvS) Schmidhuber (2019) is another important paradigm in offline RL, which eliminates Q-learning thus free of extrapolation errors. RvS learns a policy conditioned on the observed returns via supervised learning and then conditions it on a high return to generate desired behaviors Chen et al. (2021). Similar to policy constraining, RvS requires fitting the entire dataset. Therefore, the expressiveness of parameterized policies also matters in RvS. Diffusion models have the capability to represent any normalizable distribution Neal and others (2011), with the potential to effectively improve the performance of policy constraining and RvS algorithms on complex datasets.
@@ 2.2 Data Scarcity in Experience Replay
Off-policy and offline RL methods use different levels of experience replay to improve sample efficiency. Note that experience replay in some literature only refers to data reuse in off-policy RL. Here, the term broadly refers to updating the current model with rollout data from other policies. Although all previous experiences can be used for policy learning in off-policy RL, the limitation of simulation speed and the potentially huge state and action spaces may still hinder policy optimization. In offline RL, policy learning is more limited by the quality and coverage of the dataset as no further interactions are allowed. Inspired by the success of data augmentation in computer vision, some works adopt similar augmentation in RL to reduce data scarcity. RAD Laskin et al. (2020) uses image augmentation such as random cropping or rotation to improve learning efficiency in vision-based RL. Imre (2021) and Cho et al. (2022) use generative models, VAE Kingma and Welling (2013) and GAN Goodfellow et al. (2014), to augment the real dataset with generated synthetic data. However, existing works either lack fidelity when using random augmentation or are limited to simple environments due to insufficient modeling ability of particular generative models, making them difficult to be applied to more complex tasks. Diffusion models have demonstrated notable performances surpassing previous generative models in high-resolution image synthesis Ho et al. (2020). When applied to RL data, diffusion models are better suited for enhancing complex interactions.
@@ 2.3 Compounding Error in Model-based Planning
MBRL Luo et al. (2022) fits a dynamic model from online rollout data or offline datasets to facilitate decision-making. Common dynamic models mimic single-step state transitions and rewards in the dataset. When predicting with a neural dynamic model, there could be single-step errors due to the limited data support and stochastic environment transitions. Cumulative single-step errors can make planned states deviate from the dataset distribution, which causes the compounding error problem when using the model for multistep planning Xiao et al. (2019). In contrast, diffusion models with powerful modeling ability of joint distributions can operate on the trajectory level and plan for multiple time steps simultaneously, improving temporal consistency and reducing compounding errors.
@@ 2.4 Generalization in Multitask Learning
Normal RL algorithms lack generalization abilities at the task level Beck et al. (2023). Even in the same environment, changing the reward function requires retraining a policy from scratch. Existing online multitask RL Liu et al. (2021) works attempt to learn the same policy in different task environments, suffering from conflicting gradients across multiple tasks and low sample efficiency due to pure online learning. Recently, it has been a popular research direction to train a high-capacity model on multitask offline datasets and then deploy it on new tasks with or without online fine-tuning Taiga et al. (2022). Transformer-based pre-training decision models like Gato Reed et al. (2022) excel at multitask policy learning. However, they typically require high-quality datasets, large parameter sizes, and high training and inference costs. In multitask RL, designing an algorithm that can fit mixed-quality multitask datasets and generalize to new tasks emerges as a vital issue. As a powerful class of generative models, diffusion models can deal with multimodal distributions in multitask datasets, and adapt to new tasks by estimating the task distribution.
@@ 3 Foundations of Diffusion Models
We introduce the foundations of diffusion models, including two prominent formulations and several sampling techniques that are particularly important in RL tasks.
@@ 3.1 Denoising Diffusion Probabilistic Model
Assuming that the real data $x^{0}$ are sampled from an underlying distribution $q(x^{0})$, DDPM Ho et al. (2020) utilizes a parameterized diffusion process, represented as $p_{\theta}(x^{0})=\int p(x^{T})\prod_{t=1}^{T}p_{\theta}(x^{t-1}|x^{t})\mathop{}\!\mathrm{d}x^{1:T}$, to model how the pure noise $x^{T}=\mathcal{N}(\mathbf{0},\mathbf{I})$ is denoised into real data $x^{0}$. Each step of the diffusion process is represented by $x^{t}$, with $T$ indicating the total number of steps. Note that both the diffusion process and RL involve time steps; thus, we denote diffusion steps as superscripts and RL time steps as subscripts. The sequence $x^{T:0}$ is defined as a Markov chain with learned Gaussian transitions characterized by $p_{\theta}(x^{t-1}|x^{t})=\mathcal{N}(\mu_{\theta}(x^{t},t),\Sigma(x^{t},t))$. If the process is reversed as $x^{0:T}$, each step is defined by the forward transition $q(x^{t}|x^{t-1})$, which is formulated as adding Gaussian noise to the data according to a variance schedule $\beta^{1:T}$:
$x^{t}=\sqrt{\alpha^{t}}x^{t-1}+\sqrt{1-\alpha^{t}}\epsilon^{t}~{},$ (1)
where $\alpha^{t}=1-\beta^{t}$, $\epsilon^{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. From Eq. (1), we can derive a direct mapping from $x^{0}$ to $x^{t}$:
$x^{t}=\sqrt{\bar{\alpha}^{t}}x^{0}+\sqrt{1-\bar{\alpha}^{t}}{\epsilon(x^{t},t)}~{},$
where $\bar{\alpha}^{t}=\prod_{1}^{t}\alpha^{i}$. From Bayes theorem and relation between $x^{t}$ and $x^{0}$, we have
$q(x^{t-1}|x^{t},x^{0})=\mathcal{N}(\frac{1}{\sqrt{{\alpha}^{t}}}(x^{t}-\frac{\beta^{t}}{\sqrt{1-\bar{\alpha}^{t}}}\epsilon(x^{t},t)),\beta^{t}\mathbf{I})~{}.$ (2)
Eq. (2) allows us to sample $x^{T}$ from Gaussian noise and denoise step by step until we obtain $x^{0}$. However, the noise $\epsilon(x^{t},t)$ is unknown. To address this, a parameterized network $\epsilon_{\theta}$ is employed to predict the noise. Ho et al. (2020) propose the following simplified loss function for learning $\epsilon_{\theta}$, which is a weighted version of the evidence lower bound (ELBO):
$\mathcal{L}(\theta)=\mathbb{E}_{x^{0},\epsilon,t}[\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}^{t}}x^{0}+\sqrt{1-\bar{\alpha}^{t}}\epsilon,t)\|^{2}]~{},$ (3)
where $\epsilon$ is sampled from $\mathcal{N}(\mathbf{0},\mathbf{I})$.
@@ 3.2 Score-based Generative Models
Song et al. (2021) extend DDPM to continuous-time diffusion processes, and the sequence $x^{0},x^{1},\ldots,x^{T}$ is replaced with a continuous function $x^{t},t\in[0,T]$. The forwarding process can be described as a Stochastic Differential Equation (SDE):
@@
where $f(x,t)$ and $g(t)$ are pre-defined functions, and $\mathop{}\!\mathrm{d}w$ is the Brownian motion. According to Langevin dynamics, the reverse of the forwarding process is described by a reverse-time SDE:
$\mathop{}\!\mathrm{d}x=[f(x,t)-g^{2}(t)\nabla_{x}\log p_{t}(x)]\mathop{}\!\mathrm{d}t+g(t)\mathop{}\!\mathrm{d}\bar{w}~{},$
where $\bar{w}$ is the reverse Brown motion, $p_{t}(x)$ is the probability density of $x^{t}$, and $s(x)=\nabla_{x}\log p_{t}(x)$ is called the score function of $p_{t}(x)$. In practice, a parameterized score model $s_{\theta}$ is adopted to estimate the score function, which can be trained by minimizing
$\mathcal{L}(\theta)=\mathbb{E}_{x^{0},t,x^{t}}\left[\|s_{\theta}(x^{t},t)-\nabla_{x^{t}}\log p(x^{t}|x^{0})\|_{2}^{2}\right]~{}.$